OpenStudy (anonymous):

Two positive integers have a product of 200. One integer is twice the other. What are the integers?

OpenStudy (anonymous):

XY=200 2X=Y 2X^2=200 x^2=100 X=10 y=5

myininaya (myininaya):

ab=200 a=2b so 2b*b=200 b^2=100 b=-10,10 b=10 since we are doing positive numbers so a=2(10)=20

OpenStudy (anonymous):

wait i am wrong very very wrong

myininaya (myininaya):

lol

OpenStudy (anonymous):

y=20

myininaya (myininaya):

i think you divided 10 by 2 intstead of multiplied

OpenStudy (anonymous):

x=20y=10

OpenStudy (anonymous):

Lol, Thank you both so much.

myininaya (myininaya):

i gave you medal bnut